At January 1, 2020, Betty DeRose, Inc. reported work in process totaling $18,670.
This balance was made up of three jobs having the following costs:

                                    Job #171      Job #236      Job #412
  total cost at Jan. 1, 2020 .....   $4,390        $6,370        $7,910

During 2020, Betty DeRose reported the following transactions:

a. Direct materials used were: Job #171, $6,700; Job #236, $4,100; and
   Job #412, $3,900.
b. Direct labor incurred was: Job #171, $5,000; Job #236, $2,800; and
   Job #412, $3,000.
c. Overhead was applied to jobs at a rate of 85% of direct materials used.
d. Job #171, which consisted of 100 units, was completed and transferred
   to the finished goods warehouse.
e. Job #412, which consisted of 250 units, was completed and transferred
   to the finished goods warehouse.
f. 40 units from Job #171 were sold and shipped to the customer; the 40
   units were sold at a total selling price of $12,000.
g. The actual overhead cost incurred during 2020 amounted to $10,000.

Calculate the amount of gross profit reported in Betty DeRose's income
statement for 2020.
Answer #1

Job Cost Sheet

Job 171 Job 236 Job 412
Beginning balance 4390 6370 7910
Direct material 6700 4100 3900
Direct labor 5000 2800 3000
Overhead 5695 3485 3315
Total 21785 16755 18125
Unit 100
Unit price 217.85

Applied overhead = 5695+3485+3315 = 12495

Actual overhead = 10000

Over applied overhead = 2495

Adjusted cost of goods sold = (40*217.85)-2495 = 6219

Gross profit = Sales-Cost of goods sold = 12000-6219 = 5781